How many Combinations of bit strings length 9 have:                       a)    exactly three 0s? 

b)    at least seven 1s?


1
Expert's answer
2021-01-13T19:40:27-0500

If set "A" which contains "n" elements consists of "n_1" elements of the first kind, "n_2" elements of the second kind, ..., and "n_k" elements of "k"-th kind ("n=n_1+n_2+...+n_k"), the number of permutations with repetition is given by:

"\frac{n!}{n_1!\cdot n_2!\cdot ...\cdot n_k!}"


a) Let us find the number of bit strings length 9 that have exactly three 0s (and thus exactly six 1s):


"\frac{9!}{3!\cdot 6!}=\frac{9\cdot 8\cdot 7}{2\cdot 3}=84"


b) Let us find the number of bit strings length 9 that have  at least seven 1s:


"\frac{9!}{7!\cdot 2!}+\frac{9!}{8!\cdot 1!}+1=\frac{9\cdot 8}{2}+9+1=46"